OpenStudy (anonymous):

Can anyone help me solve this? ( ) = radical 5(8) + 2(20) - (8)

OpenStudy (anonymous):

ok first you have to simplify the radicals, for example (8) is going to be (4*2), since 4 is a square, you can put 2( the aquare root of 4) outside the radical and 2 inside the radical as well multiply the outside 2 with 5, and you get 10(2)

OpenStudy (anonymous):

sqrt ( 40+40-8) = sqrt(72) = sqrt(8*9) = sqrt(4*2*9) = 3*2sqrt(2) = 6sqrt2

OpenStudy (anonymous):